# Many-Body Structural Effects in Periodically Driven Quantum Batteries

### 04 Mar 2026 v1

@@ Abstract
While quantum batteries have been widely studied under static driving, their performance under periodic driving in many-body systems remains far less understood. In this Letter, we uncover structural principles establishing that many-body structure fundamentally determines the charging performance of a collective spin-1/2 quantum battery driven by a periodic Ising charger. In particular, interaction range, boundary conditions, system size, and integrability—capturing the graph connectivity, geometry, even-odd effect, and dynamics of the many-body system—emerge as critical factors in enhancing stored energy and charging power. First, we analyze how connectivity scaling and boundary geometry shape battery performance. We show that long-range interacting chargers exhibit super-extensive energy storage, approaching the well-known fundamental upper bound across broad ranges of driving periods and system sizes. In contrast, nearest-neighbor chargers achieve optimal charging only under finely tuned commensurability conditions. Moreover, we indicate that open boundary conditions (OBC) enhance robustness relative to periodic boundary conditions (PBC). Second, we further examine the impact of integrability and periodic driving on charging dynamics. We demonstrate that nonintegrability consistently enhances energy storage by suppressing conserved quantities and promoting ergodic Floquet dynamics, thereby enabling efficient population of the many-body spectrum. Through systematic structural optimization across multiple parameters, we identify long-range nonintegrability as a central resource for fast, scalable, and robust charging of collective quantum batteries. Our results elucidate how the complex structural features of many-body systems, together with periodic driving, can be harnessed to achieve efficient collective charging dynamics.
@@
Introduction.—Motivated by the fundamental limitations of classical chemical batteries and the ongoing miniaturization of quantum devices, quantum batteries (QBs) have emerged as a promising paradigm for energy storage and work extraction at microscopic scales. Since the seminal proposal by Alicki and Fannes [2], a central question has been whether quantum effects provide genuine operational advantages over classical batteries. Extensive studies have shown that uniquely quantum features, including entanglement, give rise to energy storage and release processes that outperform their classical counterparts [12, 78, 94, 21, 27, 53, 74, 34, 33, 72, 55, 48, 56, 92, 89, 84, 26, 4]. While entanglement naturally arises during the charging process and plays an important role in quantum battery (QB) [2, 8, 12], it is now understood that, especially in quantum many-body systems where “more is different,” entanglement is not the exclusive origin of quantum advantage. In fact, recent studies have demonstrated that it is neither a necessary nor a sufficient resource for enhanced charging power in many-body batteries, as maximal performance can be achieved even in the absence of strong multipartite correlations [4, 3, 41, 43, 81, 66]. This indicates that collective effects and coherence in many-body systems provide additional operational resources beyond entanglement. In particular, the emergence of collective quantum effects in interacting many-body systems can further enhance QB performance, leading to faster charging rates, increased stored energy, and greater extractable work [8, 12, 26, 4, 3, 52]. These collective advantages have motivated substantial theoretical efforts to seek optimal charging protocols and to elucidate the fundamental physical resources governing many-body QB performance [26, 4, 76, 74, 41, 52, 30, 29, 31, 80, 17, 18, 47, 49, 48, 13, 6, 77, 14, 15, 70, 58, 88, 86, 64, 59, 67, 93, 73, 5, 33]. Many of these proposals have subsequently been implemented or explored experimentally in platforms including quantum dots [54], transmons [19, 37, 28], organic semiconductors [69], superconducting circuits [36], and nuclear magnetic resonance systems [40].
@@
Collective and coherence advantages render many-body quantum batteries not only powerful platforms for energy storage but also versatile windows into complex many-body dynamics [68, 71]. However, despite these appealing features, optimizing QB performance remains highly nontrivial due to the intrinsic structural complexity of interacting many-body systems. The charging dynamics and energy landscape depend sensitively on the interaction range (including short-range, power-law, and all-to-all forms), boundary conditions (e.g., periodic or open), system size, and integrability, each of which reshapes the underlying graph connectivity and spectral structure. These intertwined structural features introduce competing length and energy scales, thereby rendering optimization a genuine multi-parameter problem. The challenge is further amplified in the presence of external driving, where the driving protocol must be optimized in conjunction with system size, interaction structure, and boundary geometry [4, 60]. While early works primarily considered quench-based unitary charging [8, 12], it has become clear that periodic and time-dependent driving provide a powerful way to engineer effective interactions, stabilize nonequilibrium steady states, and access dynamical regimes unattainable in static settings [10, 23]. In this context, Floquet or time-periodic driving has also become a powerful tool for probing nonequilibrium many-body quantum dynamics, enabling phenomena such as topological phases, prethermalization, dynamical localization, and artificial gauge fields [22, 61, 38, 57, 9, 1]. Such protocols are experimentally accessible in platforms including cold atoms in optical lattices [87, 90], trapped ions [44, 45], and superconducting qubits [91], positioning Floquet engineering as a well-established framework for introducing nonequilibrium dynamics into many-body quantum systems [24, 25, 46, 32, 62, 63, 60, 82]. Moreover, from the perspective of open quantum systems, Floquet protocols suppress decoherence and mitigate aging effects through the formation of bound states in the quasienergy spectrum of many-body systems coupled to environments [7], thereby strengthening robustness against dissipation.
@@
Combining the advantages of many-body effects with the advances in periodic driving, periodically driven many-body batteries have attracted increasing attention [60, 71, 7, 68, 20, 16]. Notably, previous studies of driven spin systems have demonstrated that long-range interactions can induce super-extensive scaling of charging power with system size, while resonance tunneling in integrable transverse-field Ising models enhances energy transfer and stabilizes stored energy [68, 60]. Nevertheless, existing works are largely confined to specific parameter regimes. For example, a recent study [71] investigated periodically kicked Ising QBs at the self-dual point with nearest-neighbor interactions in the integrable regime. Therefore, a systematic understanding of the interplay between structural effects and time-dependent driving under Floquet dynamics is crucial for advancing high-performance QBs toward practical applications.
@@
In this Letter, we investigate the structural effects in periodically driven Ising QBs within a general Floquet framework with independently tunable interaction and field durations. Our analysis transcends the kicked and self-dual limits, enabling a unified and comprehensive exploration of long-range Ising interactions with power-law decay and their nearest-neighbor counterparts under both OBC and PBC, in integrable and nonintegrable regimes. By systematically analyzing how many-body structural effects and periodic driving protocols jointly shape the charging dynamics, we characterize the stored energy and charging power across a broad parameter space. We find that periodic driving induces robust and scalable enhancements in QB performance and that, for appropriately chosen system and driving parameters, the QB attains an optimal storage energy, underscoring the central role of many-body structure in periodically driven Ising QBs. Our work provides valuable insights into the optimization of complex periodically driven many-body QBs.
@@
Short-range finite-size effects.—As a limiting case of the long-range finite-size effects, we now present the corresponding analysis for the short-range regime. At the commensurate driving period $\tau_{0}=\tau_{1}=\tau=\pi/4$, the Floquet dynamics acquire a particularly simple structure, motivating a detailed investigation of finite-size effects at this point. We therefore analyze the dependence of the stored energy on the system size at $\tau=\pi/4$. For the integrable charger under PBC, as shown in Fig. 4(a), the stored energy displays strictly periodic behavior with a period equal to the system size. A pronounced odd–even effect is observed: even system sizes achieve optimal charging, $\Delta E_{\max}=2\omega N$, at $n=N/2$, whereas odd system sizes are limited to half of the optimal value, $\Delta E_{\max}=\omega N$. In Fig. 4(b), under OBC, the dynamics remain periodic with the same system-size-dependent period; however, the maximum stored energy is uniformly bounded by $\Delta E_{\max}=\omega N$ for all system sizes, indicating that boundary effects suppress full charging even at the commensurate point.
@@
Summary.—In this Letter, we discuss how the structural properties of periodically driven Ising chargers govern the performance of QBs. Our study shows that many-body structural effects fundamentally shape both the achievable stored energy and the robustness of the charging protocol. In charging settings dominated by long-range interactions, the driving period $\tau$ acts as a crucial control parameter. At the commensurate point $\tau=\pi/2$, optimal or near-optimal charging is generally achieved, whereas periods satisfying $\tau>\pi/4$ exhibit pronounced sensitivity to boundary conditions and integrability. Furthermore, introducing unequal driving durations, $\tau_{0}\neq\tau_{1}$, reveals the robustness of nonintegrable dynamics, which sustain substantial energy storage even under temporal asymmetry, whereas integrable chargers are constrained by symmetry and exhibit reduced performance. Interestingly, the short-range one exhibits behavior fundamentally distinct from that of the long-range regime. In the nearest-neighbor case, commensurability and finite-size effects dominate the dynamics. Under PBC, integrable chargers exhibit pronounced odd-even and parity-dependent behavior, achieving full charging only for even system sizes or odd-integer couplings, whereas OBC reduces the maximal stored energy due to boundary-induced constraints. By contrast, nonintegrable dynamics smooth out these many-body structural sensitivities, enabling robust energy storage across a broad range of system sizes, boundary conditions, and coupling strengths. Overall, our results demonstrate that integrability imposes strict symmetry and parity constraints, whereas nonintegrable chaotic dynamics enhance ergodicity and reduce sensitivity to boundary effects. This interplay establishes a general framework for engineering efficient QBs, highlighting that optimal charging depends not only on interaction strength or system size but, more fundamentally, on the driving protocol and the underlying many-body structural effects. Furthermore, the maximum charging power scales approximately linearly with system size for both long-range and nearest-neighbor chargers and decreases monotonically with increasing driving period, independent of integrability or boundary conditions, as detailed in the SM [79].
@@
Acknowledgments.—Cheng Shang acknowledges the Hakubi projects of RIKEN.

### 12 Mar 2026 v2

@@ Abstract
While quantum batteries have been widely studied under static driving, their performance under periodic driving in many-body systems has received only limited attention. In this Letter, we uncover structural principles establishing that many-body structure fundamentally determines the charging performance of a collective spin-1/2 quantum battery driven by a periodic Ising charger. In particular, interaction range, boundary conditions, system size, and integrability—capturing the graph connectivity, geometry, even-odd effect, and dynamics of the many-body system—emerge as critical factors in enhancing stored energy and charging power. First, we analyze how connectivity scaling and boundary geometry shape quantum battery performance. We show that long-range interacting chargers exhibit super-extensive energy storage, approaching the well-known fundamental upper bound across broad ranges of driving periods and system sizes. In contrast, nearest-neighbor chargers achieve optimal charging only under finely tuned commensurability conditions. Moreover, we indicate that open boundary conditions (OBC) enhance robustness relative to periodic boundary conditions (PBC). Second, we further examine the impact of integrability and periodic driving on charging dynamics. We demonstrate that nonintegrability consistently enhances energy storage by suppressing conserved quantities and promoting ergodic Floquet dynamics, thereby enabling efficient population of the many-body spectrum. Through systematic structural optimization across multiple parameters, we identify long-range nonintegrability as a central resource for fast, scalable, and robust charging of collective quantum batteries. Our results elucidate how the complex structural features of many-body systems, together with periodic driving, can be harnessed to achieve efficient collective charging dynamics.
@@
Introduction.—Motivated by the fundamental limitations of classical chemical batteries and the ongoing miniaturization of quantum devices, quantum batteries (QBs) have emerged as a promising paradigm for energy storage and work extraction at microscopic scales. Since the seminal proposal by Alicki and Fannes [2], a central question has been whether quantum effects provide genuine operational advantages over classical batteries. Extensive studies have shown that uniquely quantum features, including entanglement, give rise to energy storage and release processes that outperform their classical counterparts [12, 81, 97, 21, 28, 55, 76, 35, 34, 74, 57, 49, 58, 95, 92, 87, 26, 4]. While entanglement naturally arises during the charging process and plays an important role in quantum battery (QB) [2, 8, 12], it is now understood that, especially in quantum many-body systems where “more is different,” entanglement is not the exclusive origin of quantum advantage. In fact, recent studies have demonstrated that it is neither a necessary nor a sufficient resource for enhanced charging power in many-body batteries, as maximal performance can be achieved even in the absence of strong multipartite correlations [4, 3, 42, 44, 84, 68]. This indicates that collective effects and coherence in many-body systems provide additional operational resources beyond entanglement. In particular, the emergence of collective quantum effects in interacting many-body systems can further enhance QB performance, leading to faster charging rates, increased stored energy, and greater extractable work [8, 12, 26, 4, 3, 53]. These collective advantages have motivated substantial theoretical efforts to seek optimal charging protocols and to elucidate the fundamental physical resources governing many-body QB performance [26, 4, 78, 76, 42, 53, 31, 30, 32, 83, 17, 18, 48, 50, 49, 13, 6, 79, 14, 15, 72, 60, 91, 89, 66, 61, 69, 96, 75, 5, 34]. Many of these proposals have subsequently been implemented or explored experimentally in platforms including quantum dots [56], transmons [19, 38, 29], organic semiconductors [71], superconducting circuits [37], and nuclear magnetic resonance systems [41]. In light of recent theoretical and experimental advances, QBs have become a field rich in both opportunities and challenges [27].
@@
Collective and coherence advantages render many-body QBs not only powerful platforms for energy storage but also versatile windows into complex many-body dynamics [70, 73]. However, despite these appealing features, optimizing QB performance remains highly nontrivial due to the intrinsic structural complexity of interacting many-body systems. The charging dynamics and energy landscape depend sensitively on the interaction range (including short-range, power-law, and all-to-all forms), boundary conditions (e.g., periodic or open), system size, and integrability, each of which reshapes the underlying graph connectivity and spectral structure. These intertwined structural features introduce competing length and energy scales, thereby rendering optimization a genuine multi-parameter problem. The challenge is further amplified in the presence of external driving, where the driving protocol must be optimized in conjunction with system size, interaction structure, and boundary geometry [4, 62]. While early works primarily considered quench-based unitary charging [8, 12], it has become clear that periodic and time-dependent driving provide a powerful way to engineer effective interactions, stabilize nonequilibrium steady states, and access dynamical regimes unattainable in static settings [10, 23]. In this context, Floquet or time-periodic driving has also become a powerful tool for probing nonequilibrium many-body quantum dynamics, enabling phenomena such as topological phases, prethermalization, dynamical localization, and artificial gauge fields [22, 63, 39, 59, 9, 1]. Such protocols are experimentally accessible in platforms including cold atoms in optical lattices [90, 93], trapped ions [45, 46], and superconducting qubits [94], positioning Floquet engineering as a well-established framework for introducing nonequilibrium dynamics into many-body quantum systems [24, 25, 47, 33, 64, 65, 62, 85]. Moreover, from the perspective of open quantum systems, Floquet protocols suppress decoherence and mitigate aging effects through the formation of bound states in the quasienergy spectrum of many-body systems coupled to environments [7], thereby strengthening robustness against dissipation.
@@
Combining the advantages of many-body effects with advances in periodic driving, periodically driven many-body QBs have attracted increasing attention [62, 73, 7, 70, 20, 16, 80, 54]. Periodically driven QBs were first introduced in [62]. More recently, the kicked-Ising QB model was proposed, establishing a Floquet framework to analyze the energy-injection dynamics at the self-dual point [73]. However, these prior studies are largely restricted to short-range interactions and integrable regime. Notably, recent works on driven spin systems have further shown that long-range interactions can induce super-extensive scaling of charging power with system size [80], while resonant tunneling in integrable transverse-field Ising models enhances energy transfer and stabilizes stored energy [70, 62]. These results highlight the important roles of interaction range and driving protocol in determining QB performance, yet their implications in the nonintegrable regime remain poorly understood. Therefore, a systematic understanding of how structural effects interplay with time-dependent Floquet driving remains crucial for advancing high-performance QBs.
@@
In this Letter, building upon the prior work [73, 62], we investigate the structural effects in periodically driven Ising QBs within a general Floquet framework with independently tunable interaction and field durations, covering both integrable and integrability-breaking regimes. Our analysis transcends the kicked and self-dual limits, enabling a unified and comprehensive exploration of long-range Ising interactions with power-law decay and their nearest-neighbor counterparts under both OBC and PBC, in integrable and nonintegrable regimes. By systematically analyzing how many-body structural effects and periodic driving protocols jointly shape the charging dynamics, we characterize the stored energy and charging power across a broad parameter space. We find that periodic driving robustly and scalably enhances QB performance, enabling the QB to attain optimal storage energy through appropriate choices of system and driving parameters. In particular, integrability breaking and long-range interactions emerge as key points for enhancing QB charging performance, underscoring the central role of structural effects in periodically driven Ising QBs. Our results provide valuable insights into the design and optimization of complex periodically driven many-body QBs.
@@
Short-range finite-size effects.—As a limiting case of the long-range finite-size effects, we now present the corresponding analysis for the short-range regime. At the commensurate driving period $\tau_{0}=\tau_{1}=\tau=\pi/4$, the Floquet dynamics acquire a particularly simple structure, motivating a detailed investigation of finite-size effects at this point. We therefore analyze the dependence of the stored energy on the system size at $\tau=\pi/4$. For the integrable charger under PBC, as shown in Fig. 4(a), the stored energy displays strictly periodic behavior with a period equal to the system size. A pronounced odd–even effect is observed: even system sizes achieve optimal charging, $\Delta E_{\max}=2\omega N$, at $n=N/2$, whereas odd system sizes are limited to half of the optimal value, $\Delta E_{\max}=\omega N$. In Fig. 4(b), under OBC, the dynamics remain periodic with the same system-size-dependent period; however, the maximum stored energy is uniformly bounded by $\Delta E_{\max}=\omega N$ for all system sizes, indicating that boundary effects suppress full charging even at the commensurate point. Notably, a similar analysis of the kicked Ising model shows that optimal energy storage occurs when the charger interactions act along the $x$ direction under PBC, while interactions along the $z$ direction become optimal under OBC [73]. By contrast, when the boundary conditions are interchanged—namely, $x$-direction interactions with OBC or $z$-direction interactions with PBC—the stored energy reaches only about half of the optimal value [73].
@@
Summary.— In this Letter, we discuss how the structural properties of periodically driven Ising chargers govern the performance of QBs. Building upon the kicked-Ising QB framework introduced in [73] for the integrable regime, we extend the analysis to more general periodically driven settings in the nonintegrable regime. Our study shows that many-body structural effects fundamentally shape both the achievable stored energy and the robustness of the charging protocol. In charging settings dominated by long-range interactions, the driving period $\tau$ acts as a crucial control parameter. At the commensurate point $\tau=\pi/2$, optimal or near-optimal charging is generally achieved, whereas periods satisfying $\tau>\pi/4$ exhibit pronounced sensitivity to boundary conditions and integrability. Furthermore, introducing unequal driving durations, $\tau_{0}\neq\tau_{1}$, reveals the robustness of nonintegrable dynamics, which sustain substantial energy storage even under temporal asymmetry, whereas integrable chargers are constrained by symmetry and exhibit reduced performance. Interestingly, the short-range one exhibits behavior fundamentally distinct from that of the long-range regime. In the nearest-neighbor case, commensurability and finite-size effects dominate the dynamics. Under PBC, integrable chargers exhibit pronounced odd-even and parity-dependent behavior, achieving full charging only for even system sizes or odd-integer couplings, whereas OBC reduces the maximal stored energy due to boundary-induced constraints. By contrast, nonintegrable dynamics smooth out these many-body structural sensitivities, enabling robust energy storage across a broad range of system sizes, boundary conditions, and coupling strengths. Overall, our results demonstrate that integrability imposes strict symmetry and parity constraints, whereas nonintegrable chaotic dynamics enhance ergodicity and reduce sensitivity to boundary effects. This interplay establishes a general framework for engineering efficient QBs, highlighting that optimal charging depends not only on interaction strength or system size but, more fundamentally, on the driving protocol and the underlying many-body structural effects. Furthermore, the maximum charging power scales approximately linearly with system size for both long-range and nearest-neighbor chargers and decreases monotonically with increasing driving period, independent of integrability or boundary conditions, as detailed in the SM [82].
@@
Acknowledgments.—We thank Dr. Yue Ban for valuable comments on this manuscript. We note that Ref. [73] is a prior work that introduced the kicked-Ising QB and established its basic Floquet framework at the self-dual point in the integrable regime. The present work builds on that recent study by considering a more general periodically driven setting, in which the interaction and field durations are independently tunable. Within this extended framework, we further investigate structural aspects of the charging dynamics, including long-range interactions, system-size effects, and the behavior in both integrable and nonintegrable regimes. Cheng Shang acknowledges the Hakubi projects of RIKEN.